100 ml water sample requires 6 ml of N/50 H2SO4 for neutralization upto phenolphthalein end point. Another 16 ml of the same acid was needed for further titration to methyl orange end point. Determine the type and amount of alkalinities.


1
Expert's answer
2021-02-28T06:22:21-0500

For phenolphthalein end point-

"\frac{1}{2}\ Meq\ of\ Diacidic\ base = Meq.\ of\ of \ 6ml \ \frac{N}{50}\ H2SO4"

="\frac{1\times 8}{50\times1000}=\frac{1}{6225}\ Meq\\"

"Meq.\ of\ Diacidic\ base=2\times Meq.\ of\ H2SO4 =\frac{2}{6225}\ Meq."

For methyl orange end point-

"Meq.\ of\ monoacidic\ base\ formed=Meq.\ of\ 1ml of\ extra\ H2SO4\ used"

"=\frac{1\times1}{50\times 1000}=\frac{1}{50000\ Meq.}"


Normality of diacidic alkalinity present in 100 ml water = "\frac{2\times 1000}{6225\times 100}=\frac{20}{6225}=\frac{4}{1245}\ N"